OpenStudy (anonymous):

write an explicit rule for 0.1, 0.4, 0.9, 1.6, 2.5, . . .

OpenStudy (xapproachesinfinity):

hmm my first attempt would be to write that as fractions \[\frac{1}{10},\frac{4}{10},\frac{9}{10},\frac{16}{10},\frac{25}{10}...\] observe that the top is perfect squares while the bottom stays constant